25. You have an outstanding loan with an EAR of 10.5 percent. What is the APR if interest is compounded monthly?

26. Curtis Builders is borrowing $150,000 today for 5 years. The loan is an interest-only loan with an APR of 9.5 percent. Payments are to be made annually. What is the amount of the first annual payment?

27. What is the future value of a lump sum of $100,000 invested for 6 years at an annual return of 4.0%

28. You deposit $1,000 in a savings account that pays 9 percent interest, compounded annually. How much will your account be worth in 6 years?
Answer #1

Ans 25) 11.02%

EAR = (1+r)^n - 1 Compounded Monthly EAR= EAR= (1 +10.5%/12)-12-1 11.02%

As per HomeworkLib policy one question can be answered at once.